OpenStudy (jaydah101):

please help me , i will medal and fan A line passes through the point (2, 3) and has a slope of -2. Which is the equation of the line in point-slope form? A) 2x + y = 7 B) y = -2x + 7 C) y - 3 = -2(x - 2) D) y = -12x + 5

OpenStudy (kenrick143):

Dp you know the point slope form equation?

OpenStudy (kenrick143):

*do

OpenStudy (jaydah101):

no

OpenStudy (jaydah101):

@psycho45

OpenStudy (kenrick143):

Its y-ysub1=m(x-xsub1)

OpenStudy (kenrick143):

ysub1 is 3 xsub1 is 2 and m which is the slope is -2

OpenStudy (kenrick143):

Substituting y-3=-2(x-2) simplify

OpenStudy (psycho45):

Sainte merde je déteste graphiques

OpenStudy (jaydah101):

english ?

OpenStudy (psycho45):

*sighs* I said holy crap I hate graphs

OpenStudy (psycho45):

ok if I'm not mistaking I think its B

OpenStudy (kenrick143):

Yes its B